A heating element in a stove is designed to dissipate 4,395 W when connected to 240 V.

(a) Assuming the resistance is constant, calculate the current in the heating element if it is connected to 120 V.
  A

(b) Calculate the power it dissipates at that voltage.
  W
Answer #1


power P = V^2/R

resistance R = V^2/P

R = 240^2/4395

R = 13.1 ohms

so now from ohms law

V = IR


Current I = V/R = 120/13.1 = 9.16 Amps

--------

power P = V^2/R

P = 120*120/13.1

P = 1100 watts